Fermi National Laboratory
 

Accelerator Update for November 16, 2007 - November 19, 2007

 

Friday November 16

The day shift began with Operations monitoring the Tevatron (TeV) store 5724 (20.65E30) and the Recycler stash (186.9E10), with the Antiproton Source (Pbar) stacking (50.2E10), with a beam valve (R:BV806) problem in the Recycler, with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE experiments taking beam.

At 10:06 AM, experts reported that they were certain that the Recycler beam valve had failed in a closed state. Since the stashed antiprotons would have to be aborted, Recycler experts used them in a damper study.

A MI expert began a MI-8 collimator study at 10:06 AM.

The Recycler study ended at 11:21 AM and Operations aborted the stash.

Operations began preparing MI for an access at 11:35 AM.

A TeV expert began an end-of-store study around 12:15 PM.

Vacuum technicians investigating the Recycler beam valve problem reported at 12:46 PM that a plastic airline had broken. They had completed their repairs by 1:30 PM. An RF expert also repaired MIRF01 during this access. Booster experts conducted a study during this access.

Vacuum techs came out of MI at 2:09 PM.

The TeV end-of-store study ended around 2:15 PM.

The Neutron Therapy Facility reported at 2:51 PM that they had completed treating their last patient for the day.

Operations established beam to Sci/MiniBooNE at 3:36 PM.

Pbar resumed stacking around 3:43 PM.

Operations terminated store 5724 at 4:03 PM.

At 4:37 PM, the TeV suffered a quench at sectors B0 and D0 during a TECAR power cycle.

At 5:49 PM, Operations reported that the Recycler dampers caused a large loss of antiprotons. Later, Recycler experts reverted to the old dampers.

At 5:59 PM, three Linac RF stations (LRF2, 3, & 4) tripped off. They all reset.

Operations prepared the CDF collision hall for an access at 8:13 PM. They came out about 30 minutes later.

Operations put the TeV through a dry squeeze at 8:54 PM.

At 9:25 PM, the TeV suffered a sector E3 quench. An expert came in to investigate.

Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift.

Saturday November 17

The midnight shift began with Operations monitoring the stash (111.9E10), with Pbar stacking (33.7E10), and with Sci /MiniBooNE taking beam.

At 1:56 AM, Operations reported a problem with the readbacks for the TeV sector B0 low beta magnets. Operators replaced a CAMAC card.

At 3:11 AM, Operations reported some unusual readbacks for the TeV sector B1 lead temperatures. Experts were informed.

Operations put the TeV through a dry squeeze at 4:45 AM.

Operations put the TeV through a wet squeeze at 6:20 AM.

Operations put the TeV into shot setup at 7:16 AM.

Operations reported at 7:58 AM that the Recycler had stashed two times during the midnight shift.

Operations established store 5729 at 9:33 AM with an initial luminosity of 134.17E30.

A Pbar expert began a study period at 12:53 PM.

Operations reported at 2:14 PM that the Accelerator Division Head had given NuMI permission to take beam.

Operations began turning on NuMI devices at 2:47 PM.

Operations reported at 8 PM that the Recycler had stashed three times during the last 12 hours.

Operations and experts resolved NuMI power supply, vacuum, ramping, and many other problems before finally establishing beam to NuMI at 9:54 PM.

Sunday November 18

The midnight shift began with Operations monitoring store 5729 and a stash, with Pbar stacking, and with Sci/MiniBooNE taking beam.

Operations terminated store 5729 at 12:02 AM.

Operations put the TeV into shot setup at 12:14 AM.

At 1:49 AM, Operations reported a timing problem between MI and NuMI that was tripping off the MI beam permit. Operations changed an event in the Time Line Generator and fixed the problem.

Operations established store 5730 at 2:26 AM with an initial luminosity of 123.75E30.

Operations reported at 7:58 AM that the Recycler had stashed three times during the last 12 hours.

At 2:13 PM, Operations reported a TeV separator spark at sector C49 that caused a drop in luminosity.

Operations terminated store 5730 at 5:09 PM.

Operations put the TeV into shot setup at 5:20 PM.

At 5:59 PM, all the NuMI quadrupole power supplies from areas 113 to 121 tripped off. Operations contacted experts. Later, experts reported that a capacitor had failed in each of these power supplies and would need to be replaced.

Operations established store 5731 at 7:19 PM with an initial luminosity of 135.84E30.

Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours.

Operations resumed sending beam to NuMI at 9:56 PM.

Monday November 19

The midnight shift began with Operations monitoring store 5731 and a stash, with Pbar stacking, and with NuMI and Sci/MiniBooNE taking beam.

At 4:08 AM, a NuMI kicker (I:KPS6N) tripped off and caused the NuMI beam permit to trip off. Operators investigated and found the oil sight glass for the thyratron was empty. An expert came in to investigate and reported at 7:03 AM that there was a leak in the system.

At 7:50 AM, NuMI experts reported that they would have to empty the transformer tank to repair the leak. The initial estimate was 12 hours to make repairs and then another 4-5 hours to let the oil settle.

The Plans for Monday

The plans for today are to terminate store 5731 this morning, let experts repair the TeV sector A3 thermometry crate, and then put the TeV into shot setup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/19/2007 -->    email Fermilab